OpenStudy (anonymous):

write the equation for the line that passes through points (-3,-6) and (3,2)

OpenStudy (anonymous):

y2-y1=m(x2-x1) <will give your equation m =(y2-y1)/(x2-x1)

OpenStudy (anonymous):

so then the equation would just be m=(-3+3)/(-6-2)

OpenStudy (anonymous):

this is the format (x,y)

OpenStudy (anonymous):

it doesnt matter which pair you chose but it has to remain the same.. ex m = (2+6) / (3+3)

OpenStudy (anonymous):

m is the slope of the line.

OpenStudy (anonymous):

hope that helped.. solve for m then you can write it in slope intercept form.. but im sure point slope is acceptable too ie y= mx+b (its my favorite) y=4/3x-2